OpenStudy (anonymous):

d/dx from 1 to sqrt(x) of 12t^7 dt

OpenStudy (anonymous):

what's the relation between t and x?

OpenStudy (anonymous):

or is it 12x^7?

OpenStudy (anonymous):

or sqrt(t)?

OpenStudy (anonymous):

sqrt t

OpenStudy (anonymous):

ok. in general use the following:\[\int\limits_{}^{}t^ndt=\frac{1}{t+1}t^{n+1}\]to solve these

OpenStudy (anonymous):

thus,\[12\int\limits_{1}^{\sqrt{t}}t^7dt=\frac{12}{8}((\sqrt{t})^8-\sqrt{1})\]

OpenStudy (anonymous):

forgot the exponent 8 on the sqrt(1) term but it=1 anyways

OpenStudy (anonymous):

so just simplify it and don't forget your constant of integration

OpenStudy (anonymous):

oh ic, thanks alot :)

OpenStudy (anonymous):

np

zepdrix (zepdrix):

\[\huge \frac{d}{dx}\int\limits_1^{\sqrt x} 12t^7 dt\] This looks like one of those problems where they just want you to use the FTC, Part 1 (Fundamental Theorem of Calculus). I don't think they want you to actually do the integration, since there is a d/dx in front. The limit is a sqrt x correct..? Or was it suppose to be t?

OpenStudy (anonymous):

yes sqrt x

zepdrix (zepdrix):

Here is a quick example. \[\huge \frac{d}{dx}\int\limits_0^x \cos (e^t) dt \quad = \quad \cos(e^x)\] You're integrating, but then you're differentiating, giving us back what we started with, but now in terms of a new variable, due to the limits of integration.

zepdrix (zepdrix):

In your problem, the only thing you have to be very careful about - is the fact that your upper limit is not simply "x", it's sqrt(x), so we'll have to apply the chain rule when we differentiate.

zepdrix (zepdrix):

\[\huge \frac{d}{dx}\int\limits_1^{\sqrt x}12t^7 dt\] \[\huge =12(\sqrt x)^7\frac{d}{dx}(\sqrt x)\] Understand what I did? I just plugged in the value of our upper limit for t, and then need to apply the chain rule, taking the derivative of the sqrt term.

zepdrix (zepdrix):

It's probably not what you're used to doing with integrals, so it's a little tricky to get used to :) Applying a strange rule hehe

OpenStudy (anonymous):

yes, I got it. thanks